# Distributed Thompson sampling under constrained communication

### v1

@@ D-A Simulation
$R_{B}(t)=f^{*}-\max_{\begin{subarray}{c}i\in\{1,2,\ldots M\}\\ {{x}_{t,i}\in X_{t,i}}\end{subarray}}f(\mathbf{x}_{t}^{i}),\hskip 7.11317pt\overline{R_{B}}(t)=\sum_{\tau=1}^{t}R_{B}(\tau),$
where $f^{*}=\max_{x\in\mathcal{X}}f(x).$sayInstant worst regret $R_{W}$ is the regret of the worst performing agent at time step $t,$ and we define its sum to be the quantity $\overline{R_{W}}:$
$R_{W}(t)=f^{*}-\min_{i\in\{1,2,\ldots M\}}f(x_{t,i}),\hskip 7.11317pt\overline{R_{W}}(t)=\sum_{\tau=1}^{t}R_{W}(\tau).$
[…]
Instant average regret $R_{A}$ is the average of the regrets of all agents at time step $t,$ and we define its sum to be $\overline{R_{A}}:$
[…]
D-B Discussion
@@
Our Distributed Thompson Sampling algorithm was able to achieve the extrema of the Ackley and Rosenbrock objective functions, and thus is effective at the Bayesian optimization task. Our theoretical result bounds expectation of Bayesian Simple Regret, $R_{SB}(t),$ with the bound dependent on the size of the largest complete subgraph of the communication network between agents. In the star graph configuration, the largest complete subgraph contained $2$ vertices for any star graph containing at least $2$ vertices. Interestingly, in Figure 1, we see similar performance in the $3,5,$ and $7$ agent cases, with significantly worse performance in the single agent case. This matches intuition motivated by our theoretical result, as the $3,5,$ and $7$ agent cases all have largest complete subgraph size of $2,$ as opposed to the single agent case, where the largest complete subgraph is of size $1$
E Conclusion
@@
In this paper, we proposed a Distributed Thompson Sampling algorithm to address the multi-agent Bayesian optimization problem under constrained communication. We develop a bound on Bayesian Simple Regret for this approach, where the bound is dependent on properties of the largest complete subgraph of the graph encoding communication structure between agents. With our bound, we show that in multi-agent communication networks, Bayesian Simple Regret will converge faster than in the sequential case (with the same number of rounds). Additionally, we demonstrate the efficacy of our algorithm with regret analysis on optimization test functions, illustrating faster convergence with graphs possessing larger connected subgraphs. Future work will focus on giving a tighter regret bound, and developing adjustments for the Distributed Thompson Sampling algorithm to better control the exploration/exploitation tradeoff.

### v3

@@ IV-A Simulation
We also define the Instant simple regret $R_{S}$, and its sum $\overline{R_{S}}$, as follows:
$R_{S}(t)=f^{*}-\max_{\begin{subarray}{c}i\in\{1,2,\ldots M\},\\ {\tau\in\{1,2,\ldots,t\}}\end{subarray}}f(x_{t,i}),\hskip 7.11317pt\overline{R_{S}}(t)=\sum_{\tau=1}^{t}R_{S}(\tau),$
where $f^{*}=\max_{x\in\mathcal{X}}f(x).$ Numerical results were constructed in a Python implementation built upon the BOTorch package [31]. The code used to generate the simulation and corresponding Figure 1 is available at https://github.com/sabzer/distributed-bo. The Gaussian processes utilized the Matérn kernel with parameter $\nu=\frac{5}{2}.$ The numerical simulations were run over $T=50$ timesteps. Simulations were run based on two test functions for the objective function: Ackley, which has many local maxima and one global minima in the origin, and Rosenbrock, which contains a large valley in which the global minima is situated. The equations of the aforementioned objective functions and their plots are available in Appendix C. Since we were solving a maximization problem, we multiplied the canonical definitions of these functions by $-1$ for the purpose of our simulation. For the communication networks in our simulations, we used Erdős-Rényi random graphs of $20$ agents with connectivities of $0.2,$ $0.4,$ and $0.6$ [32]. The connection probabilities are the probability that each edge from the complete graph of $20$ agents appears in the corresponding random graph.
IV-B Discussion
@@
Our theoretical result bounds Bayesian average regret, $R_{AB}(t),$ and Bayesian simple regret, $R_{SB}(t),$ with the bound dependent on the structure of the communication network between agents. Our distributed Thompson sampling algorithm was able to achieve the extrema of the Ackley and Rosenbrock objective functions in numerical implementation, and thus is effective at the Bayesian optimization task. Our theoretical results suggests that the distributed Thompson sampling algorithm implementation favors highly connected communication graphs. This is apparent from a lower Bayesian average regret bound when the communication graph can be decomposed into a few large disjoint complete subgraphs, and a lower Bayesian simple regret bound when the largest complete subgraph of the communication graph has a larger number of agents. Our numerical results support this intuition, for in Figure 1, we see better regret convergence for Erdős-Rényi graphs of higher connectivity. This result holds for both Ackley and Rosenbrock objective functions, and for both Instant simple and average regret.
V Conclusion
@@
In this paper, we proposed a distributed Thompson sampling algorithm to address the multi-agent Bayesian optimization problem under constrained communication. We develop bounds on Bayesian average regret and Bayesian simple regret for this approach, where the bound is dependent on properties of the largest complete subgraph of the graph encoding communication structure between agents. With our bound, we show that in connected multi-agent communication networks, both Bayesian average regret and Bayesian simple regret will converge faster with distributed Thompson sampling than in the sequential single-agent case, with the same number of rounds. Additionally, we demonstrate the efficacy of our algorithm with regret analysis on optimization test functions, illustrating faster convergence with well connected communication graphs. Future work will focus on developing a tighter regret bound, and further tailoring the distributed Thompson sampling algorithm towards the constrained communication case by leveraging the data communicated between agents.